(a) Find and identify the traces of the quadric surface

x2 + y2z2 = 4, given the plane.


x = k

y=k

z=k

(IDENTIFY TRACE AND SHAPE OF THE TRACE)
Answer #1

x2 + y2z2 = 4

x = k ==> y2z2 = 4-k2 ==> hyperbola

y = k ==> x2z2 = 4-k2 ==> hyperbola

z = k ==> x2 + y2 = 4+k2 ==> circle.

Given equation is a one sheet hyperboloid.